# DarkDNS: Revisiting the Value of Rapid Zone Update

### v1

@@ Abstract.
Malicious actors exploit the DNS namespace to launch spam campaigns, phishing attacks, malware, and other harmful activities. Combating these threats requires visibility into domain existence, ownership and nameservice activity that the DNS protocol does not itself provide. To facilitate visibility and security-related study of the expanding gTLD namespace, ICANN introduced the Centralized Zone Data Service (CZDS) that shares daily zone file snapshots of new gTLD zones. However, a remarkably high concentration of malicious activity is associated with domains that do not live long enough make it into these daily snapshots. Using public and private sources of newly observed domains to identify this activity, we discover that even with the best available data there is a considerable visibility gap. We find that the daily snapshots miss at least 1% of newly registered and short-lived domains, which are almost always registered with malicious intent. In reducing this critical visibility gap using public sources of data, we demonstrate how more timely access to TLD zone changes can help better prevent abuse. We hope that this work sparks a discussion in the community on how to effectively and safely revive the concept of sharing Rapid Zone Updates for security research.
journalyear: 2024conference: arXiv pre-print; submitted; under peer-reviewbooktitle: arXiv pre-print - Under peer-review
@@ 1. Introduction
Malicious actors exploit (abuse) the DNS namespace to launch spam campaigns, phishing attacks, malware, and other harmful activities. In many dimensions the DNS ecosystem is more opaque than other aspects of Internet transport. Unlike BGP, DNS is a pull protocol, so learning internal dynamics requires an entry point i.e., a domain name. Consequently, in absence of knowing this exact entry point any abuse that lies behind a domain will remain opaque to everyone except the targets.
@@
This inherent opacity has motivated many attempts to improve visibility into the DNS ecosystem to facilitate research, analysis, and improved operational security. A prominent example is the ICANN Centralized Zone Data Service (CZDS) which coordinates collection and restricted sharing of daily zone file snapshots of participating TLDs. However, previous literature, in attempting to provide independent analysis of vulnerabilities and exploits, has found that the 24-hour zone files, while an unprecedented enabler of DNS-related security research, leave vital visibility gaps in studies of abuse and exploits (Akiwate et al., 2020; Akiwate et al., 2021; Security and Stability Advisory Committee, 2024). Since so much malicious activity often begins soon after a domain is registered for that purpose, and the CZDS snapshots are updated only daily, commercial threat intelligence services (Domain Tools, 2024) have tried to fill the gap by detecting domains soon after registration.
@@
In this work we try to quantify, using public and private sources of data, the gaps in the current models of data sharing, both regulatory and market-driven. Of particular interest are domains that are created and removed in under 24 hours, and thus do not appear in the CZDS zone files. The predominant reason for such transient domains is a fraudulent or malicious activity. For example, if registrars receive evidence of malicious use or flag potentially fraudulent payment methods after they process the registration, the registrar will remove the domain from the zone immediately i.e., before the next CZDS zone snapshot. Unfortunately, abuse prevention currently relies solely on the registrars. Each registrar must independently identify the same signals to block the same threat actors, who continuously move across registrars to evade detection. This results in an endless cat-and-mouse game where attackers maintain the upper hand. Fine-grained visibility into TLD zone changes can enable verified third parties such security researchers to build signals to prevent abuse and raise the cost for threat actors in this ecosystem.
@@
In using public and private sources of data to get fine grained visibility into TLD zone changes, we find that we reduce but do not completely close the significant visibility gap. Using ground truth from a medium-size European ccTLD registry, we find that we are able to detect only one-third of these transient domains even with the best public available data. Given the measurable expansion of cybercrime, and the failure of current data sharing models to effectively combat it, security-conscious TLDs should consider offering a subscription service to a feed of changes to zones immediately as they occur. Verisign offered such a service for the .com TLD 15 years ago (ICANN, 2007) but later terminated it due to its potential for abuse. We recognize these concerns of abuse but believe now is the time to develop a framework to manage these concerns, just as ICANN is doing for access to even more sensitive types of registration data (ICANN, 2023).
@@
    We provide an open data feed of newly registered domains (including transient) we discover.
@@ 2. Related Work
Barron et al.  (Barron et al., 2019) found significant correlation between early domain name deletions and potentially malicious activities, and noted that often zone files do not reveal whether a domain is registered. Other studies have found that spammer domains generally had short lifetimes, and that blocklists generally did not detect them until they were actively used and reported rather than when initially registered (Hao et al., 2013).
Affinito et al.  (Antonia Affinito and Raffaele Sommese and Gautam Akiwate and Stefan Savage and Kimberly Claffy and Geoffrey M. Voelker and Alessio Botta and Mattijs Jonker, 2022) found patterns of suspiciously short lifetimes domain names in some TLDs (e.g., 80% under .xyz), many of which were reported as malicious. Their study revealed that domains registered under some TLDs (e.g., .xyz, .icu) are taken down within a few days, while others (e.g., .top) remain in the zone longer before removal.
Commercial threat intelligence services have tried to narrow this visbility gap by monitoring DNS queries in traffic. In 2018 Domain Tools reported that most newly observed domains become inactive after 4 hours (Pawel Foremski and Paul Vixie, 2018), presumably because they have been detected and flagged as malicious and no longer serve their purpose. Observing DNS queries from their CDN, Akamai identified 20.1% (13M) of all newly registered domain they detected as malicious. (Akamai, 2024). In 2022, Palo Alto Newtorks reported that 70% of newly registered domains detected from passive DNS data and zone files across 1530 TLDs were malicious or suspicious or not safe for work (Chen and Wang, 2022).
@@
A recent study is relevant to our goal of detection of newly observed domains without acces to proprietary data. Last year, Sommese et al. . demonstrated that CT transparency logs are as a reliable and up-to-date source of domain names for addressing the visibility gaps in ccTLDs, most of which do not participate in CZDS. They were able to use CT logs to independently reconstructure more than half of the domains in several ccTLDs for which they had the complete zone file. Their work focused on the spatial coverage; we apply CT logs to analyze temporal gaps, i.e., how quickly we can detect new and transient domains (Sommese et al., 2023).
@@ 3. Methodology
Our approach to estimating the extent of newly registered domains not captured by daily zone snapshots relies on intersecting public data sources: available zone snapshots, certificate transparency logs (Laurie et al., 2013), active DNS resolution measurements, and registration data (RDAP) (Newton and Hollenbeck, 2015). Our pipeline has five steps: identify domains for which certificates are issued but are not present in the latest CZDS snapshots; collect RDAP data; monitor DNS changes of newly observed domains by performing active measurements; cross-validate inference with RDAP data; and identify transient (short-lived) domains. We feed the results of each measurement into Kafka topics and store them in Parquet format in our object storage for longitudinal analysis.
@@
We start with the daily zone files extracted by our collector, which is populated with all latest zones snapshots available from ICANN CZDS. We use the open-source Certstream (Calidog, 2024) to captures logs of newly issued certificates 111 We only consider PreCertificate entries, because they must always be published before the issuance of the final certificate (Laurie et al., 2013), and extract domain names from the Common Name (CN) and the Subject Alternative Name (SAN) fields of these certificates. We discard domain names that already appeared in the latest daily zone snapshots. This step has tree limitations. First, some CAs do not always perform Domain Validation step [^2] Per Subsection 4.2.1 of the CA Browser Forum baseline requirements (Forum, 2024), a CA may reuse cached validation information collected no more than 398 days prior to issuing the Certificate. (Ma et al., 2023), so domains with certificates less than 398 days old may no longer exist (§4). Second, zone file publication may be delayed by days, leading to inaccurate inference of domain existence. Lastly, we are able to detect domains only if a related certificate is issued. This step yields a stream of potentially newly registered domain names that we feed into our analysis pipeline.
@@
We verify how accurately we detect a newly registered domain using its RDAP-reported registration date and time. We deploy a script in Azure cloud that retrieves newly registered domains from our Kafka topic and uses the whoisit Python library to perform RDAP queries for each domain. We leverage Azure functions to avoid aggressive rate-limiting by cycling measurements over different IPv4 addresses. To minimize overhead, we did not respond to query failures with retry mechanisms.
@@
We check that the RDAP registration timestamp is consistent (within 24 hours) with the observed certificate timestamp. The difference between these two timestamps provides a ground truth indication of how correctly and promptly we detect newly registered domain names. This step only works when our RDAP resolution succeeded ($\approx 97\%$ of cases. [^3] We missed $\approx 3\%$ due to RDAP rate limiting or other collection errors.)
@@ 4. Results
TLD Category Dec Nov Jan Total Zone NRD Coverage NRD (%)
com 1 109 80411098041\,109\,8041 109 804 1 127 72711277271\,127\,7271 127 727 1 505 04415050441\,505\,0441 505 044 3 742 57537425753\,742\,5753 742 575 8 467 64184676418\,467\,6418 467 641 44.2%
xyz 87 0518705187\,05187 051 114 582114582114\,582114 582 107 740107740107\,740107 740 309 373309373309\,373309 373 649 010649010649\,010649 010 47.7%
shop 99 6609966099\,66099 660 76 6267662676\,62676 626 107 675107675107\,675107 675 283 961283961283\,961283 961 775 253775253775\,253775 253 36.6%
online 76 6937669376\,69376 693 76 6747667476\,67476 674 109 964109964109\,964109 964 263 331263331263\,331263 331 648 922648922648\,922648 922 40.6%
bond 81 2658126581\,26581 265 75 7797577975\,77975 779 84 9978499784\,99784 997 242 041242041242\,041242 041 292 552292552292\,552292 552 82.7%
top 74 1347413474\,13474 134 82 7468274682\,74682 746 83 8378383783\,83783 837 240 717240717240\,717240 717 532 363532363532\,363532 363 45.2%
net 71 9227192271\,92271 922 79 6607966079\,66079 660 84 3208432084\,32084 320 235 902235902235\,902235 902 643 030643030643\,030643 030 36.7%
org 53 7675376753\,76753 767 53 3775337753\,37753 377 76 4007640076\,40076 400 183 544183544183\,544183 544 481 870481870481\,870481 870 38.1%
site 47 8794787947\,87947 879 46 6954669546\,69546 695 65 8016580165\,80165 801 160 375160375160\,375160 375 465 542465542465\,542465 542 34.4%
store 38 6993869938\,69938 699 42 9314293142\,93142 931 50 2795027950\,27950 279 131 909131909131\,909131 909 326 383326383326\,383326 383 40.4%
Others 333 000333000333\,000333 000 328 570328570328\,570328 570 380 551380551380\,551380 551 1 042 12110421211\,042\,1211 042 121 3 009 57530095753\,009\,5753 009 575 34.6%
Total 2 073 87420738742\,073\,8742 073 874 2 105 36721053672\,105\,3672 105 367 2 656 60826566082\,656\,6082 656 608 6 835 84968358496\,835\,8496 835 849 16 292 1411629214116\,292\,14116 292 141 42.0%
Table 1. Newly Registered Domain names per month per Top 10 TLDs (in numbers of newly registered domains).
Recall, our methodology uses CT logs to identify use of domains before they show up in the CZDS zone snapshots. From 1 Nov 2023 to 31 Jan 2024, we found 6.8 million domain names (Table 1) that are not in the CZDS zone snapshots. These $6.8M$ domain names are either transient domains or eventually show up in the next CZDS zone snapshots. Most of these domains (3.7M) were in .com, followed by .xyz (300K) and .shop (284K).
When comparing our list of detected newly registered domains to the diff between two daily CZDS zone snapshots, we found that our methodology identified 42% of the newly registered domains before they were published in the zone snapshots in our 3 month observation period. The remaining 58% did not have certificates issued immediately. Since our goal is to identify transient domains i.e., , domains that do not show up in zone snapshots, this comparison is purely functional.
@@ 4.1. Detection Speed
Figure 1 plots the time difference between when we recorded the domain (i.e., fetched it from CT Logs [^4] We use the Certstream-reported timestamp because neither precertificates nor CT logs provide an insert timestamp) and the RDAP-reported timestamp. We detected half of the newly registered domains within the first 45 minutes of their existence, and $\approx$30within the first 15 minutes. The distribution has a long tail toward 1 day (less than 2% have a difference greater than a day), which we believe derives from our misclassification of newly registered domains (e.g., due to incorrect extraction of the second-level domain (SLD) using the Public Suffix List) or zones published with several days of delay. The time differences across TLDs can be explained by the operational update times of their zones. For example, the .com and .net registries update these TLD zones on average every 60 seconds while other gTLDs registries update their zones every 15-30 minutes. We validated this assumption by probing the zones of Figure 1 for SOA serial changes, and found consistent timestamps.

### v2

@@ Abstract.
Malicious actors exploit the DNS namespace to launch spam campaigns, phishing attacks, malware, and other harmful activities. Combating these threats requires visibility into domain existence, ownership and nameservice activity that the DNS protocol does not itself provide. To facilitate visibility and security-related study of the expanding gTLD namespace, ICANN introduced the Centralized Zone Data Service (CZDS) that shares daily zone file snapshots of new gTLD zones. However, a remarkably high concentration of malicious activity is associated with domains that do not live long enough make it into these daily snapshots. Using public and private sources of newly observed domains, we discover that even with the best available data there is a considerable visibility gap in detecting short-lived domains. We find that the daily snapshots miss at least 1% of newly registered and short-lived domains, which are frequently registered with likely malicious intent. In reducing this critical visibility gap using public sources of data, we demonstrate how more timely access to TLD zone changes can provide valuable data to better prevent abuse. We hope that this work sparks a discussion in the community on how to effectively and safely revive the concept of sharing Rapid Zone Updates for security research. Finally, as a contribution of this work, we are releasing a public live feed of newly registered domains, with the aim of enabling further research in early abuse identification.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM Internet Measurement Conference; November 4–6, 2024; Madrid, Spainbooktitle: Proceedings of the 2024 ACM Internet Measurement Conference (IMC ’24), November 4–6, 2024, Madrid, Spaindoi: 10.1145/3646547.3689021isbn: 979-8-4007-0592-2/24/11ccs: Networks Naming and addressingccs: Networks Network measurement
@@ 1. Introduction
Malicious actors exploit (abuse) the DNS namespace to launch spam campaigns, phishing attacks, malware, and other harmful activities. In many dimensions the DNS ecosystem is more opaque than other aspects of Internet transport. Unlike BGP, DNS is a pull protocol, so learning internal dynamics requires an entry point i.e., a domain name. Without knowing this entry point, any abuse that lies behind a domain remains opaque to everyone except the targets.
@@
This inherent opacity has motivated many attempts to improve visibility into the DNS ecosystem to facilitate research, analysis, and improved operational security. A prominent example is the ICANN Centralized Zone Data Service (CZDS) which coordinates collection and restricted sharing of daily zone file snapshots of participating TLDs. However, previous studies have found that the 24-hour zone files, while an unprecedented enabler of DNS-related security research, leave vital visibility gaps in studies of abuse and exploits (Akiwate et al., 2020; Akiwate et al., 2021; Security and Stability Advisory Committee, 2024). Since so much malicious activity often begins soon after a domain is registered for that purpose, and the CZDS snapshots are updated only daily, commercial threat intelligence services (Domain Tools, 2024) have tried to fill the gap by detecting domains soon after registration.
@@
In this work we try to quantify, using public and private sources of data, the gaps in the current models of data sharing, both regulatory and market-driven. We focus on domains that are created and removed in under 24 hours, and thus do not appear in the CZDS zone files. The likely predominant reason for such transient domains is a fraudulent or malicious activity. For example, if registrars receive evidence of malicious use or flag potentially fraudulent payment methods after they process the registration, the registrar will remove the domain from the zone immediately i.e., before the next CZDS zone snapshot. Unfortunately, abuse prevention currently relies solely on the registrars. Each registrar must independently identify the same signals to block the same threat actors, who continuously move across registrars to evade detection. This results in an endless cat-and-mouse game where attackers maintain the upper hand. Fine-grained visibility into TLD zone changes can enable verified third parties such as security researchers to build signals to prevent abuse and raise the cost for threat actors.
@@
In using public and private sources of data to get fine grained visibility into TLD zone changes, we can reduce but not completely close the significant visibility gap. Using ground truth from a medium-size European ccTLD registry, we detected only one-third of these transient domains even with the best public available data. Given the measurable expansion of cybercrime, and the failure of current data sharing models to effectively combat it, security-conscious TLDs should consider offering a subscription service to a feed of changes to zones immediately as they occur. Verisign offered such a service for the .com TLD 15 years ago (ICANN, 2007) but later terminated it due to its potential for abuse. We recognize these concerns of abuse but believe now is the time to develop a framework to manage these concerns, just as ICANN is doing for access to even more sensitive types of registration data (ICANN, 2023).
@@
    We provide an open data feed of newly registered domains (including transient) we discover (OpenINTEL, 2024).
  3. (3)
    We investigate the infrastructural landscape of transient domains and the possible reasons for their early removal.
@@ 2. Related Work
Barron et al.  (Barron et al., 2019) found significant correlation between early domain name deletions and potentially malicious activities, and noted that often zone files do not reveal whether a domain is registered. Other studies have found that spammer domains generally had short lifetimes, and that blocklists generally did not detect them until they were actively used and reported rather than when initially registered (Hao et al., 2013). Affinito et al.  (Antonia Affinito and Raffaele Sommese and Gautam Akiwate and Stefan Savage and Kimberly Claffy and Geoffrey M. Voelker and Alessio Botta and Mattijs Jonker, 2022) found patterns of suspiciously short lifetimes domain names in some TLDs (e.g., 80% under .xyz), many of which were reported as malicious. Their study revealed that domains registered under some TLDs (e.g., .xyz, .icu) are taken down within a few days, while others (e.g., .top) remain in the zone longer before removal.
Commercial threat intelligence services have tried to narrow this visibility gap by monitoring DNS queries in traffic. In 2018 Domain Tools reported that most newly observed domains become inactive after 4 hours (Pawel Foremski and Paul Vixie, 2018), presumably because they have been detected and flagged as malicious and no longer serve their purpose. Observing DNS queries from their CDN, Akamai identified 20.1% (13M) of all newly registered domain they detected as malicious. (Akamai, 2024). In 2022, Palo Alto Newtorks reported that 70% of newly registered domains detected from passive DNS data and zone files across 1530 TLDs were malicious or suspicious or not safe for work (Chen and Wang, 2022).
@@
A recent study is relevant to our goal of detection of newly observed domains without access to proprietary data. Last year, Sommese et al.  (Sommese et al., 2024). demonstrated that CT transparency logs are a reliable and up-to-date source of domain names for addressing the visibility gaps in ccTLDs, most of which do not participate in CZDS. They used CT logs to independently reconstruct more than half of the domains in several ccTLDs for which they had the full zone file as ground truth. Their work focused on spatial coverage whereas we use CT logs to analyze temporal gaps, i.e., how quickly we detect new and transient domains (Sommese et al., 2024).
@@ 3. Methodology
Our approach to estimating the extent of newly registered domains not captured by daily zone snapshots relies on public data sources: available zone snapshots, certificate transparency logs (Laurie et al., 2013), active DNS resolution measurements, and registration data (RDAP) (Newton and Hollenbeck, 2015). Unlike passive DNS, our approach does not require the domain to be actively queried by users, but the the domain must have an issued certificate. Our pipeline has five steps: identify registered (pay-level) domains for which certificates are issued but are not present in the latest CZDS snapshots; collect RDAP data; monitor DNS changes of newly observed domains by performing active measurements; cross-validate inference with RDAP data; and identify transient (short-lived) domains. We feed the results of each measurement into Kafka topics and store them in Parquet format in our object storage for longitudinal analysis.
@@
We start with the daily zone files extracted by our collector, which is populated with all latest zones snapshots available from ICANN CZDS. We use the open-source Certstream (Calidog, 2024) to captures logs of newly issued certificates, 111 We only consider PreCertificate entries, because they must always be published before the issuance of the final certificate (Laurie et al., 2013) and extract domain names from the Common Name (CN) and the Subject Alternative Name (SAN) fields of these certificates. We discard domain names that already appeared in the latest daily zone snapshots. This step has three limitations. First, some CAs do not always perform Domain Validation step [^2] Per Section 4.2.1 of the CA Browser Forum baseline requirements (Forum, 2024), a CA may reuse cached validation information collected no more than 398 days prior to issuing the Certificate. (Ma et al., 2023), so domains with certificates fewer than 398 days old may no longer exist (§4). Second, zone file publication may be delayed by days, leading to inaccurate inference of domain existence. Lastly, we are able to detect domains only if a related certificate is issued. This step yields a stream of potentially newly registered domain names that we feed into our analysis pipeline. While, this stream contains only domains with an associated issued certificate, it represents a publicly accessible data feed for researchers, in contrast with commercial passive DNS feeds. We release this feed publicly at (OpenINTEL, 2024).
@@
We verify how accurately we detect a newly registered domain using its RDAP-reported registration date and time. In addition, we collect registrar and registrant information, the latter of which is typically redacted to protect privacy. To collect RDAP data, we deploy a script in Azure cloud that retrieves newly registered domains from our Kafka topic and uses the whoisit Python library to perform RDAP queries for each domain. We leverage Azure functions to avoid aggressive rate-limiting by cycling measurements over different IPv4 addresses. To minimize overhead, we did not retry failed queries.
@@
We check that the RDAP registration timestamp is consistent (within 24 hours) with the observed certificate timestamp. The difference between these two timestamps provides a ground truth indication of how correctly and promptly we detect newly registered domain names. This step only works when our RDAP resolution succeeded ($\approx 97\%$ of cases [^3] We missed $\approx 3\%$ due to RDAP rate limiting or other collection errors.).
@@ 4. Results
Table 1. Top 10 TLDs ranked by count of newly registered domains (NRDs) for Nov 23 - Jan 24). The Zone NRD columns shows how many NRDs appeared in the daily zone file snapshots.
TLD Category Nov Dec Jan Total Zone NRD Coverage NRD (%)
com 1 127 72711277271\,127\,7271 127 727 1 109 80411098041\,109\,8041 109 804 1 505 04415050441\,505\,0441 505 044 3 742 57537425753\,742\,5753 742 575 8 467 64184676418\,467\,6418 467 641 44.2%
xyz 114 582114582114\,582114 582 87 0518705187\,05187 051 107 740107740107\,740107 740 309 373309373309\,373309 373 649 010649010649\,010649 010 47.7%
shop 76 6267662676\,62676 626 99 6609966099\,66099 660 107 675107675107\,675107 675 283 961283961283\,961283 961 775 253775253775\,253775 253 36.6%
online 76 6747667476\,67476 674 76 6937669376\,69376 693 109 964109964109\,964109 964 263 331263331263\,331263 331 648 922648922648\,922648 922 40.6%
bond 75 7797577975\,77975 779 81 2658126581\,26581 265 84 9978499784\,99784 997 242 041242041242\,041242 041 292 552292552292\,552292 552 82.7%
top 82 7468274682\,74682 746 74 1347413474\,13474 134 83 8378383783\,83783 837 240 717240717240\,717240 717 532 363532363532\,363532 363 45.2%
net 79 6607966079\,66079 660 71 9227192271\,92271 922 84 3208432084\,32084 320 235 902235902235\,902235 902 643 030643030643\,030643 030 36.7%
org 53 3775337753\,37753 377 53 7675376753\,76753 767 76 4007640076\,40076 400 183 544183544183\,544183 544 481 870481870481\,870481 870 38.1%
site 46 6954669546\,69546 695 47 8794787947\,87947 879 65 8016580165\,80165 801 160 375160375160\,375160 375 465 542465542465\,542465 542 34.4%
store 42 9314293142\,93142 931 38 6993869938\,69938 699 50 2795027950\,27950 279 131 909131909131\,909131 909 326 383326383326\,383326 383 40.4%
Others 328 570328570328\,570328 570 333 000333000333\,000333 000 380 551380551380\,551380 551 1 042 12110421211\,042\,1211 042 121 3 009 57530095753\,009\,5753 009 575 34.6%
Total 2 105 36721053672\,105\,3672 105 367 2 073 87420738742\,073\,8742 073 874 2 656 60826566082\,656\,6082 656 608 6 835 84968358496\,835\,8496 835 849 16 292 1411629214116\,292\,14116 292 141 42.0%
Recall, our methodology uses CT logs to identify use of domains before they show up in the CZDS zone snapshots. From 1 Nov 2023 to 31 Jan 2024, we found 6.8 million domain names (Table 1) that did not appear in the CZDS zone snapshot before we saw them in the CT logs. These $6.8M$ domain names are either transient domains or eventually show up in the next CZDS zone snapshots. Most of these domains (3.7M) were in .com, followed by .xyz (300K) and .shop (284K).
When comparing our list of detected newly registered domains to the diff between two daily CZDS zone snapshots, we found that our methodology identified 42% of the newly registered domains before they were published in the zone snapshots in our 3-month observation period. The remaining 58% did not have certificates issued immediately. Since our goal is to identify transient domains i.e., , domains that do not show up in zone snapshots, this comparison is purely functional.
@@ 4.1. Detection Speed
Figure 1 plots the time difference between when we recorded the domain (i.e., fetched it from CT Logs [^4] We use the Certstream-reported timestamp because neither precertificates nor CT logs provide an insert timestamp) and the RDAP-reported timestamp. We detected half of the newly registered domains within the first 45 minutes of their existence, and $\approx$30within the first 15 minutes. The distribution has a long tail toward 1 day (less than 2% have a difference greater than a day), which we believe derives from our misclassification of newly registered domains (e.g., due to incorrect extraction of the second-level domain (SLD) using the Public Suffix List) or zones published with several days of delay. The time differences across TLDs can be explained by the operational update times of their zones (i.e., the domain needs to be active in the zone before a certificate can be issued). For example, the .com and .net registries update these TLD zones on average every 60 seconds while other gTLDs registries update their zones every 15-30 minutes. We validated this assumption by probing the zones of Figure 1 for SOA serial changes, and found consistent timestamps.